OpenStudy (anonymous):

What formula do I use to solve this, I just need to know the formula only:) You would like to have $1000000 in 38 years by making regular deposits at the end of each month in to an annuity that pays 4% compounded monthly. Determine the amount of each monthly deposit. What formula?

OpenStudy (anonymous):

is it... A=P(1+rt) or P(1+r/m)^mt=R[(1+r/m)^mt-1)/(r/m)] A=R[(1+r/m)^mt-1)/(r/m)]

OpenStudy (anonymous):

or A=P(1+r/m)mt

jimthompson5910 (jim_thompson5910):

You use the formula A = R[ (1+r/m)^(mt)-1 )/(r/m)]

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

You're welcome You can find more info here http://en.wikipedia.org/wiki/Annuity_%28finance_theory%29